Grade upto college level Physical Chemistry

Estimate the average S-F bond energy in SF6. The values of standard enthalpy of formation of SF6(g), S(g)and F(g) are : - 1100,275 and 80 kJ mol-1 respectively.

Shane Macguire
12 Years agoGrade upto college level

2 Answers

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
Given S(s) + 3F2(g) → SF6(g); ∆ H = - 1100 kJ ….(i)
(s) → S(g); ∆ H = 275 kJ …(ii)
1/2 F2(g) → F(g); ∆ H = 80 kJ …(iii)
To get SF6(g) → S(g) + 6F(g) we can proceed as
(ii) + 6 * (iii) - (i)
∴ SF6(g) → S(g) + 6F(g); ∆ H = 1855 kJ
Thus average bond energy for S-F bond = 1855/6 = 309.16kJ

Thanks
jitender
askIITians Faculty
Kushagra Madhukar
5 Years ago
Dear student,
Please find the answer to your question
 
Given S(s) + 3F2(g) → SF6(g); ∆ H =  – 1100 kJ ….(i)
(s) → S(g); ∆ H = 275 kJ …(ii)
1/2 F2(g) → F(g); ∆ H = 80 kJ …(iii)
To get SF6(g) → S(g) + 6F(g) we can proceed as
(ii) + 6 * (iii) - (i)
∴ SF6(g) → S(g) + 6F(g); ∆ H = 1855 kJ
Thus average bond energy for S-F bond = 1855/6 = 309.16kJ
 
Thanks and regards,
Kushagra